OpenStudy (anonymous):

Solve sin^2(θ) = 3cos^2(θ) What I have gotten so far is: sin^2/cos^2=3 tan^2=3 x= arctan(sqrt(3)) x=pi/3 I know tan(2pi)=0 so I have the solution of pi/3+2piN. This is reflected in multiple of the choices. How do I get the rest of the equations? Also, one of the possible answers is "not one of these choices"

myininaya (myininaya):

ISe the identity that sin^2(theta)=1-cos^2(theta) do not do any divisions as you have done above because cos can be 0 sometimes

myininaya (myininaya):

USE *

OpenStudy (triciaal):

solve aren't you just suppose to find the angle? tan^2 theta = 3 so tan theta = sqrt 3 is x the radius of the circle? are you finding the equation of the circle with that angle?

OpenStudy (anonymous):

The possible answers are : a: pi/3, 5pi/3 b: pi/3+2pin, 2pi/3+2pin, 4pi/3+2pin, 5pi/3+2pin c: pi/3+2pin,5pi/3 + 2pin d: none of these

OpenStudy (triciaal):

cos theta = 1/2, -1/2

myininaya (myininaya):

@sccitesla As I was saying you do not won't to divide the initial equation by cos because cos can sometimes be 0 (and you do not divide by 0) If you do, this could eliminate some answers that you should have had

myininaya (myininaya):

Use the identity mentioned above 1-cos^2(theta)=sin^2(theta)

OpenStudy (anonymous):

How can i use that density with no stray additions. Am i supposed to add them in?

myininaya (myininaya):

I see sin^2(theta) so I would replace sin^2(theta) with 1-cos^2(theta) like so: \[1-\cos^2(\theta)=3\cos^2(\theta) \]

myininaya (myininaya):

can you solve this?

OpenStudy (anonymous):

...sorry brain fart.

OpenStudy (anonymous):

im still in the same place i was before i believe...... though hopefully i missed something

OpenStudy (anonymous):

arccos(.5)=pi/3

OpenStudy (anonymous):

+-.5=cos(x) cos(2pi/3)= -.5 cos(2pi/3)= -.5 cos(4pi/3)= -.5 cos(5pi/3)= .5

OpenStudy (anonymous):

no answers have been eliminated

myininaya (myininaya):

I'm going to add cos^2(theta) on both sides giving me \[1=4 \cos^2(\theta)\] divide both sides by 4 \[\frac{1}{4}=\cos^2(\theta)\] Now to get rid of the square you must square root both sides this will give you two equations to solve.

myininaya (myininaya):

Oh I think you got it right there 1/2=cos(x) or -1/2=cos(x)

myininaya (myininaya):

@sccitesla You found your answer.

myininaya (myininaya):

You have already stated it.

OpenStudy (anonymous):

Those are not in the choices

OpenStudy (anonymous):

nvm

myininaya (myininaya):

so you don't see pi/3, 2pi/3, 4pi/3, 5pi/3 all of those +2npi

OpenStudy (anonymous):

i see. Stupid freaking things like this is what gets me stuck -_-

OpenStudy (anonymous):

wait... one second

OpenStudy (anonymous):

so the only answer that has solutions in both positive and negative is b. so that must be the answer. Ok thanks